OpenStudy (anonymous):

What is the slope in the equation: y = 6x - 300?

OpenStudy (anonymous):

$$y=mx \pm c$$ hence slope here is 6

OpenStudy (anonymous):

6 is the answer y=mx+c , where m=6

OpenStudy (muhammad_nauman_umair):

As 6x -y -300=0 And we know that Slope for a line is {-(a/b)} where a=6 and b=-1 then slope will be6